OpenStudy (anonymous):

√14700=√100x147=10√147 What is the mistake?

hartnn (hartnn):

no mistake,its correct :)

OpenStudy (anonymous):

The paper says to find the student's mistake/:

hartnn (hartnn):

maybe middle step should be written as: \[\sqrt{100X147}\] but final answer is correct

jimthompson5910 (jim_thompson5910):

The step \[\Large \sqrt{100}\sqrt{147}\] is missing, but it's not 100% critical to have (esp if you know what's going on)

OpenStudy (cruffo):

147 = 3*49 So the student hasn't finished "simplifying"

OpenStudy (cruffo):

\[\sqrt{14700} = \sqrt{100 \times 147} = 10\sqrt{147} = 10\sqrt{49 \times3} = 10\times 7 \sqrt{3} = 70\sqrt{3}\]

OpenStudy (anonymous):

Thank you all!